OpenStudy (aaronandyson):

A = 3i-4j ,B= -i-4j find:

OpenStudy (aaronandyson):

\[A+B , |A+B|\]

OpenStudy (jtug6):

So apparently from what you've drawn it's wanting vector addition of vectors a and b, and then the magnitude of the sum of vectors a and b?

OpenStudy (aaronandyson):

Yes!

OpenStudy (jtug6):

i'm still relatively new to vectors so i may be wrong but for the first one it's as simple as addingthe components of vects A/B. so add your coefficients of i and j,

OpenStudy (aaronandyson):

2i-8j

OpenStudy (jtug6):

i'd say thats correct =). of course someones probably going to want to check this just to make sure haha. okay and the second one youre going to want to find the magnitude of vect |A + B| so in order to find the magnitude you need to take the sqrt of the sum of each component squared.

OpenStudy (aaronandyson):

What? Do I square A and B and then add then?

OpenStudy (jtug6):

mmmmm youre basically squaring the coefficients of i/j of both a/b and then summing all of that. then taking the square root of all of that. i'll help you get started. sec.

OpenStudy (aaronandyson):

-i = 1 right?

OpenStudy (jtug6):

mmmm -i's coefficient would be -1 since -1 * i = -i but the negatives dont matter anyway since it's being squared

OpenStudy (aaronandyson):

so, -1^2 +(-4)^2?

OpenStudy (jtug6):

ok no actually i take that back sorry it's actually the magnitude of our new vector |A+B| that we found for part A! we aren't using the same vector components. so ignore what i drew actually. again sorry! we're going to use that 2i -8j vector we found =)

OpenStudy (jtug6):

so we're squaring 2 and -8, summing them, and then taking the square root of all of that.

OpenStudy (aaronandyson):

60?

OpenStudy (jtug6):

er 64 + 4 right? cause 8^2 = 64 and 2^2 = 4

OpenStudy (jtug6):

that'd be 68 =]

OpenStudy (jtug6):

and then square root of that.

OpenStudy (aaronandyson):

8.2?

OpenStudy (jtug6):

yes! if they want you to round to a certain decimal amount then be careful but otherwise yes 8.2 :P

OpenStudy (aaronandyson):

THANKS!

OpenStudy (jtug6):

were these multiple choice or?

OpenStudy (jtug6):

and np but still i'd be careful i'm really new to vectors still ;p

OpenStudy (jtug6):

@agent0smith am i right? :3

OpenStudy (agent0smith):

A = 3i-4j ,B= -i-4j \[\large A+B = 2i - 8j\] \[\Large |A + B| = \sqrt{2^2+8^2}\]

OpenStudy (jtug6):

thats what i had yayyyy

rebeccaxhawaii (rebeccaxhawaii):

MY DANNO ;u;

OpenStudy (agent0smith):

I just didn't feel like reading it

OpenStudy (jtug6):

;l all good.